3 of us at Grosmont today - Angie, Chris H and me.
Given that we were limited in the amount of work that only three of us could undertake we concentrated on preparing the K1 for its boiler insurance cold exam (arranged for next Friday).
We completed cleaning the mud hole doors and washout plugs, thoroughly cleaned the fire box and smoke box and removed fire hole bottom protector plate and safety valves.
Towards the end of the day we removed a valve ring which had been stuck in the back LH valve head and started cleaning both heads on the same valve.
All three of our loco's at Grosmont are still to be moved into Deviation Shed. This did not hamper our progress today and as it worked out, it is a good job the Q6 was still outside as it will be working tomorrow (Repton failed with a blown piston packing today and Eric was trapped behind a damaged point).
We knew that the Q6 regulator gland had been leaking by when it was used on Wednesday so we repacked the gland this afternoon.
The next Grosmont working party will be on Wednesday then both days next weekend. It is imperative that we push on with the winter maintenance on all of our loco's at Grosmont so please come and join us if you can.